OpenStudy (knogirl):

can someone please explain : 3+ sqrt(18) / 1 + sqrt (8)

OpenStudy (christopher1448):

3+ sqrt(18) / 1 + sqrt (8)Primarily, a sqrt simplification drill must begin with "sqrt" Any other syntax is not allowed. For example, expressions like sin^(-1)(sqrt(300)) are not accepted. Supported are, for example: Square roots of integers which are > 1 : sqrt(1024) Square roots of integers followed by variables (Which may be raised to a power): sqrt(288x^3y^-5zq) Square roots of variables (which may be raised to a power): sqrt(zx4y5)

OpenStudy (christopher1448):

is this helpful?

OpenStudy (ashauntay):

√8 = √(4×2) = √4 × √2 = 2√2

OpenStudy (christopher1448):

^

OpenStudy (christopher1448):

@Hermes

OpenStudy (christopher1448):

....

OpenStudy (christopher1448):

)_)

OpenStudy (knogirl):

\[3+\sqrt{18}\div 1+\sqrt{8}\]

OpenStudy (knogirl):

I don't understand how to solve

OpenStudy (knogirl):

I understand that \[\sqrt{18}= 3\sqrt{2} and \sqrt{8} = 2\sqrt{2}\]